A 25.0 mL solution of Ba(OH)₂ is neutralized with 31.5 mL of 0.250 M HBr. What is the concentration of the original Ba(OH)₂ solution?


1
Expert's answer
2021-05-10T01:25:47-0400


m(solution)=31,5ml/0.25=126g

M(сoncentration)=V(sub)/V(sol)

n(Ba(OH)2)=0.025l/22.4l/mol=0,001116mol

m(Ba(OH)2)=171g/mol*0,001116mol=0.19

M(Ba(OH)2)=0.19g/126g=0.015